Use technology to solve the following problem: A bottler of drinking water fills plastic bottles with a mean volume of 1,002 milliliters (mL) and standard deviation  The fill volumes are normally distributed. What proportion of bottles have volumes less than 


1
Expert's answer
2021-07-21T12:30:08-0400

"\mu=1002"

"\sigma=7"

"P(x<1000)-?"


"z=\frac{x-\mu}{\sigma}=\frac{1000-1002}{7}=-0.29"


"P(x<1000)=P(z<-0.29)=0.38591"